#16 value investing beats growth investing over extended periods of time. A Corporate Bond-Face Value Par Value of $1000 with an 8% coupon interest rate. This Bond has no call features. Fill in the table below: Bought at Par or Face Value. What is the purchase price? 65. Bought at a 20% discount to Par or Face Value, what is the purchase price? Bought at a 50% discount to Par or Face Value. What is the purchase yield? Bought at an 80% discount to Par or Face Value, what is the current bond yield? #17 Fundamental Analysis is an investors best friend.
Answer #1

1) If baought at par or face value then current rpcie will be = $ 1000

-------------------------------------------------------------------------------------

2) 20% is calculated on par or face value. So purchase price = (1-20%) x 1000 = $800

------------------------------------------------------------------------------------

3) Annual Coupon 8% or 1000 x 8% = $80

Purchase Price = (1-50%) x 1000 = 500

So Yield = 80 /500 = 16%

----------------------------------------------------------------------------

4) Annual coupon = 8% x 1000 = 80

Purchase price = (1-80%) x 1000 = 200

Yield = 80/200 = 40 %

-------------------------------------------------

Note : Assume yield formula of current yield. Price = (1-Disc rate) x Par value.